Using FOIL, simplify the expression "(3x + 2)(3x - 2)." Show that a particular factoring formula leads to the same answer. I know how to use FOIL, I am not sure of another way to solve this other than foil.
Answer by richwmiller(17219)


This is a formula for factoring x^2-y2
x^2-y^2=(x+y)*(x-y)
unfactoring
(x+y)*(x-y)=x^2-y^2
Here is how it applies here.
9x^2-4
BTW
Here is an alternative to FOIL
Break up the first factor into two parts 3x and 2
3x*(3x-2)+2*(3x-2)=9x^2-6x+6x-4=9x^2-4